OpenStudy (anonymous):

Please help! dont know how to go about solving this definite integral. -2 to 4 x^2(x^3+8)dx

OpenStudy (zarkon):

$$\int_{-2}^{4}x^2(x^3+8)dx$$ $$u=x^3+8$$ $$du=3x^2dx$$ $$\frac{du}{3}=x^2dx$$ $$\frac{1}{3}\int_{0}^{72}udu$$ $$=\frac{1}{3}\left.\frac{u^2}{2}\right|_{0}^{72}=864$$